An octagon has eight interior angles, five of which sum to 845 degrees. Two of the remaining three angles are supplementary and two of the remaining three angles are complementary. Find the angle measures of the remaining three angles.

Answer by cleomenius(959)

The total of the 8 angles for an octagon is (n - 2) * 180 degrees = 1080 degrees.
1080- 845 = 235 degrees.
x + y + z = 235
235 = 180 = 55 degrees, this one of the angles.
90 degrees - 55 degrees = 35 degrees, this another angle.
180 degrees -35 degrees = 145 degrees, this is the third angle.
Cleomenius.
55 + 35 + 145 + 845 = 1085.
This does check.
Cleomenius